# Parameter Efficient Fine-tuning via Cross Block Orchestration for Segment Anything Model

### 28 Nov 2023 v1

@@ 1 Introduction
In the field of computer vision, the majority of leading PEFT methodologies have focused on image classification tasks [16, 39, 15]. These approaches demonstrate that large classification models can be fine-tuned by injecting a small number of parameters. As evidence, their performance can match or even surpass that of vanilla full fine-tuning [15]. Influenced by the success of the existing PEFT methods, one can directly apply them in fine-tuning segmentation models [25, 38], e.g., Segment Anything Model (SAM) [20]. However, fine-tuning segmentation models often requires a heavier adjustment of parameters, which poses a challenge for current PEFT methods that are limited by the Hidden Markov Chain (HMC) along layers [32], as this characteristic limits them to making substantial adjustments of the projection directions in the entire parameter space for new scenarios, as shown in Fig. 1.
In this paper, we equip PEFT with a CrOss-BLock OchesTration mechanism to enable the adaptation of SAM to various downstream scenarios, dubbed as SAM-COBOT. The goal of SAM-COBOT is to explicitly integrate cross-block orchestration to enhance the flexibility and reliability of adjusting projection directions. Specifically, we first propose an inter-block communication (IBC) module, which introduces a learnable relation matrix to capture interdependence and facilitate communication among different blocks. The communication is realized by adjusting the coefficient set of each PEFT block’s parameter space. We treat all the coefficient sets of PEFT block’s parameter space as a tensor, and use the learnable relation matrix to capture cross-slice information for adjusting each coefficient set. Consequently, IBC allows projection directions to influence each other in the entire parameter space. Subsequently, we introduce an intra-block enhancement (IBE) module, which includes a linear project head whose weights are generated from a hyper-complex layer, to ensure that any coordinated adjustments made to the projection directions achieve a greater impact on the entire parameter space.
@@ 3.1 Hyper-complex Number
In the hyper-complex number, $a_{0}$ is the real part, $a_{1}{j_{1}}+a_{2}{j_{2}}+\cdots+a_{n-1}{j_{n-1}}$ is the imaginary part. For simplicity, we consider $n=4$ in this work. (Details about $8$-dimensional and $16$-dimensional hyper-complex numbers are given in Supplementary Material.) The imaginary part of a $4$-dimensional hyper-complex number satisfies:
@@
The geometric interpretation of $j_{1}$, $j_{2}$, and $j_{3}$ can be understood as rotations in $\mathbb{R}^{3}$ space: the $j_{1}$ rotation represents the rotation from the X-axis to the Y-axis in the plane intersecting both axes, the $j_{2}$ rotation represents the rotation from the Z-axis to the X-axis in the plane intersecting with only the Z-axis, and finally, the $j_{3}$ rotation represents a rotation from Y-axis to Z-axis in a plane that intersects both axes. The negative counterparts $-j_{1}$, $-j_{2}$, and $-j_{3}$ represent reverse rotations of their respective positive counterparts. It is worth noting that, unlike the real and complex numbers, multiplication of the imaginary part of the 4-dimensional hyper-complex number is not commutative, for example, $j_{1}j_{2}=j_{3}$ and $j_{2}j_{1}=-j_{3}$.
@@ 3.3 Parameter Efficient Fine-tuning (PEFT)
Drawbacks. While existing PEFT methods have shown promising results when adapted to SAM, we observe a fundamental conflict in this framework. Given that fine-tuning segmentation models typically necessitates a heavier adjustment of parameters to align projection directions in the parameter space for new scenarios compared to classification models. However, these PEFT methods introduce only a limited number of parameters in each layer, which can only make relatively small adjustments of projection directions due to the limitation of Hidden Markov Chain (HMC) along layers. To solve this, we focus on achieving proper adjustments by mitigating the constraint of the HMC, while introducing nearly zero newly trainable parameters.
@@ 4.2 Inter-block Communication
After that, we propose to introduce a learnable relation matrix for achieving inter-block communication. To achieve this, we need to conceptualize all coefficient set matrices in the entire parameter space as a single tensor. Denote the number of transformer blocks in SAM’s image encoder as $L$, and each block contains a PEFT module with a coefficient set that we proposed. We treat the diagonal matrix of each coefficient set as an individual slice in this tensor $\mathcal{T}$, which is derived as:
@@
According to the characteristics of gradient propagation in deep learning, each frontal slice $\mathbf{\Lambda}_{i}\in\mathbb{R}^{V\times V}$ of the tensor $\mathcal{T}\in\mathbb{R}^{V\times V\times L}$ is updated sequentially, and update the tensor $\mathcal{T}$ is slow. To avoid the cross-frontal-slice information loss in the tensor $\mathcal{T}$ during learning, we introduce the idea of a special tensor product, i.e., T-product.
@@ 5.1 Experimental Setup
(a) Initialization
(b) TRCAN [12]
(c) NWPU [9, 8, 7]
(d) ADMOE [26]
Figure 5: Visualization of Relation Matrix (RM) on three datasets. Initialization: we initialize the RM as a diagonal matrix.
@@
Implementation Details. In all of our experiments, we employ the ViT-Base version of SAM [20] as our backbone, integrating a box prompt for its prompt encoder input. In line with previous studies [25, 38], we apply a random perturbation to each bounding box, varying between 0 and 50 pixels. Our training employs the Adam optimizer [19]. For medical image segmentation, the initial learning rate is set to $1.25\times 10^{-6}$, and the weight decay is $5\times 10^{-4}$ with one image per mini-batch. The number of fine-tuning epochs is set to 25. For natural and remote sensing image segmentation, we follow SonarSAM [38], the initial learning rate is set to $10^{-4}$, and the weight decay is $5\times 10^{-5}$ with one image per mini-batch. The number of fine-tuning epochs is set to 20.
CoS RM HL ADMOE NWPU TRCAN
@@
Strategy of RM ADMOE NWPU TRCAN
@@
Linear HL ADMOE NWPU TRCAN
@@ 5.2 Ablative Studies
Ablation of Main Components. Here, we do an ablation study to show the benefit brought by each component of our proposed SAM-COBOT, i.e., coefficient set (CoS), relation matrix (RM), and hyper-complex layer (HL) on three datasets, including ADMOE [26], NWPU [9, 8, 7] and TRCAN datasets [12]. We use Adaptformer [6] as the baseline here and show its results in row 1. As shown in Table  1, comparing row 2 to row 1, we can see slight performance gains brought by the coefficient set, as it introduces more parameter space to be optimized. Furthermore, the results are further boosted by large margins after introducing the relation matrix. This reveals that the relation matrix controls dependencies among different layers. Finally, by integrating the hyper-complex layer, the results reveal clear performance gains, e.g., 1.2% on ADMOE dataset, demonstrating the effective interaction within each layer.
@@
Effects of Relation Matrix (RM). In Table 2, we demonstrate the efficacy of our proposed learnable relation matrix by comparing it with a fixed matrix initialized with random values. We can observe a significant performance improvement with our method, e.g., 1.1% DSC on ADMOE dataset [26]. Beside, Fig. 5 shows the different distribution w.r.t. various scenarios. The aboved result suggests that our relation matrix captures valuable cross-block information.
@@
Linear Layer or Hyper-complex Layer. Table 3 presents a comparison between our proposed hyper-complex layer and a standard linear layer, which is a commonly used module for communicating among channels, i.e., projection directions. The results reveal improvements in performance, e.g., 0.4% DSC ADMOE dataset [26]. This suggests that the orthogonality facilitated by the hyper-complex layer plays a beneficial role in enhancing intra-layer communication.
@@ 5.3 Main Results
Comparing to SOTA. We compare our approach against several prevailing PEFT techniques for SAM, including LoRA [14] and Adapterformer [6], on 10 datasets across three domains in the computer vision community. We present their original results and also show our results (by plugging SAM-COBOT in these methods) in Table 4. As shown in Table 4, our SAM-COBOT becomes the new state-of-the-art. Remarkably, our SAM-COBOT, with minimal parameter overhead, significantly enhances both LoRA and Adaptformer. This improvement is particularly noticeable in medical image segmentation. By way of illustration, our SAM-COBOT boosts Adaptformer by 1.2% and 1.0% in terms of mIoU on ADMOE [26] and SEGRAP [35] dataset, respectively. Notably, our method achieves 0.5% DSC gains on LoRA on BRAST dataset, despite starting from a lower performance (i.e., 84.8%) than the baseline (i.e., 85.1%). Notably, LoRA achieves a 1.1% improvement through lightweight fine-tuning on average by incorporating 147.4K parameters. In contrast, SAM-COBOT, with a mere addition of 0.9K parameters, achieves an additional 0.7% enhancement in performance. The above results underscore the robustness and generalization capabilities of our SAM-COBOT.
@@
Figure 7: Qualitative segmentation results on three scenarios, i.e., (a) natural image segmentation on COCO dataset [22], (b) natural image segmentation on TRCAN [12] dataset, (c) remote sensing image segmentation on SSDD [46] dataset, (d) remote sensing image segmentation on NWPU [9, 8, 7] dataset and (e) medical image segmentation on ADMOE [26] dataset. “Lightweight”: freezes all the backbone parameters and only tunes SAM’s lightweight mask decoder.
@@
Different Backbones. We extend our fine-tuning to include larger-scale backbones, e.g., ViT-Base and ViT-Large, reinforcing the versatility of our method. This is evaluated on the SSDD and ADMOE datasets, where, as Table 5 illustrates, performance improvements are observed consistently. These results demonstrate the generalization of our approach across various transformer architectures.
Different Hidden Dimensions. In Fig. 6, we compare our method with a baseline model, specifically Adaptformer, across various dimensions of the hidden space $V$. Overall, our method demonstrates distinct advantages in all dimensional settings. It is noteworthy that at lower dimensions, e.g., $V\leq 16$, our method achieves more pronounced performance improvements, exceeding 1.6% on the ADMOE dataset [26]. This observation underscores the efficiency of our method in facilitating interaction among bases, particularly when the number of bases, or $V$, is constrained.

### 28 Mar 2024 v2

@@ 1 Introduction
In the field of computer vision, the majority of leading PEFT methodologies have focused on image classification tasks [16, 40, 15]. These approaches demonstrate that large classification models can be fine-tuned by injecting a small number of parameters. As evidence, their performance can match or even surpass that of vanilla full fine-tuning [15]. Influenced by the success of the existing PEFT methods, one can directly apply them in fine-tuning segmentation models [26, 39], e.g., Segment Anything Model (SAM) [20]. However, fine-tuning segmentation models often necessitates a heavier adjustment of parameters as the output space for segmentation is often much larger and more varied compared to classification, which poses a challenge for current PEFT methods that are limited by the Hidden Markov Chain (HMC) along layers [41, 21]. The memoryless nature of HMC implies that each layer’s state is influenced only by its adjacent layers, thus readily leading to minor adjustments of the projection directions in the entire parameter space for new scenarios, as shown in Fig. 1.
In this paper, we equip PEFT with a CrOss-BLock OrchesTration mechanism to enable the adaptation of SAM to various downstream scenarios, dubbed as SAM-COBOT. The goal of SAM-COBOT is to explicitly integrate cross-block orchestration to enhance the flexibility and reliability of adjusting projection directions. Specifically, we first propose an inter-block communication (IBC) module, which introduces a learnable relation matrix to capture interdependence and facilitate communication among different blocks. The communication is realized by adjusting the coefficient set of each PEFT block’s parameter space. We treat all the coefficient sets of PEFT block’s parameter space as a tensor, and use the learnable relation matrix to capture cross-slice information for adjusting each coefficient set. Consequently, IBC allows projection directions to influence each other in the entire parameter space. Subsequently, we introduce an intra-block enhancement (IBE) module, which includes a linear project head whose weights are generated from a hyper-complex layer, to ensure that any coordinated adjustments made to the projection directions achieve a greater impact on the entire parameter space.
@@ 3.1 Hyper-complex Number
In the hyper-complex number, $a_{0}$ is the real part, $a_{1}{j_{1}}+a_{2}{j_{2}}+\cdots+a_{n-1}{j_{n-1}}$ is the imaginary part. For simplicity, we consider $n=4$ in this work. The imaginary part of a $4$-dimensional hyper-complex number satisfies:
@@
The geometric interpretation of $j_{1}$, $j_{2}$, and $j_{3}$ can be understood as rotations in $\mathbb{R}^{3}$ space: the $j_{1}$ rotation represents the rotation from the X-axis to the Y-axis in the plane intersecting both axes, the $j_{2}$ rotation represents the rotation from the Z-axis to the X-axis in the plane intersecting both axes, and finally, the $j_{3}$ rotation represents a rotation from Y-axis to Z-axis in a plane that intersects both axes. The negative counterparts $-j_{1}$, $-j_{2}$, and $-j_{3}$ represent reverse rotations of their respective positive counterparts. It is worth noting that, unlike the real and complex numbers, multiplication of the imaginary part of the 4-dimensional hyper-complex number is not commutative, for example, $j_{1}j_{2}=j_{3}$ and $j_{2}j_{1}=-j_{3}$.
@@ 3.3 Parameter Efficient Fine-tuning (PEFT)
Drawbacks. Although existing PEFT methods can be directly integrated with SAM, we notice an opening question in this intuitive solution. Fine-tuning segmentation models often necessitates a heavier adjustment of parameters to align projection directions in the parameter space for new scenarios compared to classification models. However, these PEFT methods introduce only a limited number of parameters in each layer, which can only make relatively small adjustments of projection directions due to the limitation of Hidden Markov Chain (HMC) along SAM’s layers. Although some methods, e.g., LST [38], seem to bypass this issue by integrating a learnable side adapter, the updating of each layer in the side adapter is also limited to interactions with its adjacent layers. Consequently, the limitations of HMC still exist. In contrast, we devise a plug-and-play method which directly mitigates the constraint of the HMC for existing PEFT methods while introducing nearly zero training efforts.
@@ 4.2 Inter-block Communication
After that, we propose to introduce a learnable relation matrix for achieving inter-block communication. To do this, we need to conceptualize all coefficient set matrices in the entire parameter space as a single tensor. Denote the number of transformer blocks in SAM’s image encoder as $L$, and each block contains a PEFT module with a coefficient set that we proposed. We treat the diagonal matrix of each coefficient set as an individual slice in this tensor $\mathcal{T}$, which is derived as:
@@
Then, according to the characteristics of gradient propagation in deep learning theory, i.e., chain rule, each frontal slice $\mathbf{\Lambda}_{i}\in\mathbb{R}^{V\times V}$ of the tensor $\mathcal{T}\in\mathbb{R}^{V\times V\times L}$ is updated sequentially, and thus update the tensor $\mathcal{T}$ is often slow. To avoid the cross-frontal-slice information loss in the tensor $\mathcal{T}$ during learning, we introduce the idea of a special tensor product, i.e., T-product.
@@ 5.1 Experimental Setup
Implementation Details. In all of our experiments, we employ the ViT-Base version of SAM [20] as our backbone, integrating a box prompt for its prompt encoder input. In line with previous studies [26, 39], we apply a random perturbation to each bounding box, varying between 0 and 50 pixels. Our training employs the Adam optimizer [19]. For medical image segmentation, the initial learning rate is set to $1.25\times 10^{-6}$, and the weight decay is $5\times 10^{-4}$ with one image per mini-batch. The number of fine-tuning epochs is set to 25. For natural and remote sensing image segmentation, we follow SonarSAM [39], the initial learning rate is set to $10^{-4}$, and the weight decay is $5\times 10^{-5}$ with one image per mini-batch. The number of fine-tuning epochs is set to 20. More details are provided in the supplementary material.
CoS RM HL ADOME NWPU TRCAN
@@
Strategy of RM ADOME NWPU TRCAN
@@
Linear HL ADOME NWPU TRCAN
@@ 5.2 Ablative Studies
Ablation of Main Components. Here, we do an ablation study to show the benefit brought by each component of our proposed SAM-COBOT, i.e., coefficient set (CoS), relation matrix (RM), and hyper-complex layer (HL) on three datasets, including ADOME [27], NWPU [9, 8, 7] and TRCAN datasets [12]. We use Adaptformer [6] as the baseline in row 1 of Table  1. Comparing row 2 to row 1, we can see slight performance gains brought by the coefficient set, as it introduces more parameter space to be optimized. Then, solely introducing the hyper-complex layer shows limited improvement as it can only adjust projection directions within each layer. Moreover, the results are further boosted by large margins after introducing the relation matrix, showing its capability to capture interdependencies among different layers. Finally, by integrating the hyper-complex layer, the results reveal clear performance gains, e.g., 1.2% on ADOME dataset.
@@
Effects of Relation Matrix (RM). In Table 2, we demonstrate the efficacy of our proposed learnable relation matrix by comparing it with a fixed matrix initialized with random values. We can observe a significant performance improvement with our method, e.g., 1.1% DSC on ADOME dataset [27].
@@
Linear Layer or Hyper-complex Layer. Table 3 presents a comparison between our proposed hyper-complex layer and a standard linear layer, which is a commonly used module for communicating among channels, i.e., projection directions. The results reveal improvements in performance, e.g., 0.4% DSC ADOME dataset [27]. This suggests that the orthogonality facilitated by the hyper-complex layer plays a beneficial role in enhancing intra-layer communication.
@@ 5.3 Main Results
Comparing to SOTA. We compare our approach against several prevailing PEFT techniques for SAM, including LoRA [14] and Adapterformer [6], on 10 datasets across three domains in the computer vision community. We present their original results and also show our results (by plugging SAM-COBOT in these methods) in Table 4. As shown in Table 4, our SAM-COBOT becomes the new state-of-the-art. Remarkably, our SAM-COBOT, with minimal parameter overhead, significantly enhances both LoRA and Adaptformer. This improvement is particularly noticeable in medical image segmentation. By way of illustration, our SAM-COBOT boosts Adaptformer by 1.2% and 1.0% in terms of mIoU on ADOME [27] and SEGRAP [35] dataset, respectively. Notably, our method achieves 0.5% DSC gains on LoRA on BRAST dataset, despite starting from a lower performance (i.e., 84.8%) than the baseline (i.e., 85.1%). Notably, LoRA achieves a 1.1% improvement through lightweight fine-tuning on average by incorporating 147.4K parameters. In contrast, SAM-COBOT, with a mere addition of 0.9K parameters, achieves an additional 0.7% enhancement in performance. The above results underscore the robustness and generalization capabilities of our SAM-COBOT. More results are provided in the supplementary material.
@@
Figure 6: Qualitative segmentation results on three scenarios, i.e., (a) natural image segmentation on COCO dataset [23], (b) natural image segmentation on TRCAN [12] dataset, (c) remote sensing image segmentation on SSDD [48] dataset, (d) remote sensing image segmentation on NWPU [9, 8, 7] dataset and (e) medical image segmentation on ADOME [27] dataset. “Lightweight”: freezes all the backbone parameters and only tunes SAM’s lightweight mask decoder.
@@
Different Backbones. We extend our fine-tuning to include larger-scale backbones, e.g., ViT-Base and ViT-Large, reinforcing the versatility of our method. This is evaluated on the SSDD and ADOME datasets, where, as Table 5 illustrates, performance improvements are observed consistently. These results demonstrate the generalization of our approach across various transformer architectures.
Different Hidden Dimensions. In Fig. 5, we compare our method with a baseline model, specifically Adaptformer, across various dimensions of the hidden space $V$. Overall, our method demonstrates distinct advantages in all dimensional settings. It is noteworthy that at lower dimensions, e.g., $V\leq 16$, our method achieves more pronounced performance improvements, exceeding 1.6% on the ADOME dataset [27]. This observation underscores the efficiency of our method in facilitating interaction among bases, particularly when the number of bases, or $V$, is constrained.
@@ 6 Conclusion
7 Acknowledgments
This work was supported by NSFC 62322604, 62176159, Natural Science Foundation of Shanghai 21ZR1432200, and Shanghai Municipal Science and Technology Major Project 2021SHZDZX0102.